# The Critical LYZ equation in Kähler Geometry

### v2

@@ 2.1. The $k$-Hessian operators and the special Lagrangian operator
Lemma 2.2 (Yuan [44], Wang–Yuan [42]).
@@ 4.1. Viscosity solutions
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain. Suppose $w$ is an (n1)(n-1)-admissible function and a viscosity subsolution of (45) in $\Omega$, and $v$ is a viscosity solution of (45) in $\Omega$. If $w\leq v$ on $\partial\Omega$, then $w\leq v$ in $\Omega$.
@@
Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible. As $v$ is a viscosity solution (and hence a supersolution), by the definition of the supersolution, we have
@@ 4.2. Weak solutions in the sense of currents
    The function $\frac{1}{2}v^{2}$ is a subsolution to the same equation.
@@
Let $v$ be a weak solution (respectively, subsolution) in the sense of currents to the equation $\sigma_{n}+\sigma_{n-1}=0$. Then $v$ is also a viscosity solution (respectively, subsolution).
@@
The proof for supersolutions follows similarly, using the corresponding definition of viscosity supersolutions. ∎
@@ 4.3. A Liouville theorem to the equation $\sigma_{n}+\sigma_{n-1}=0$ in $\mathbb{C}^{n}$
Let $v:\mathbb{C}^{n}\rightarrow\mathbb{R}$ be a $C^{1}$ function that is (n1)(n-1)-subharmonic and such that $v+|z|^{2}$ is plurisubharmonic. Assume there exists a positive constant $C$ such that $\|v\|_{C^{1}(\mathbb{C}^{n})}\leq C$ and $\Delta v\leq C$ in the weak sense (see Remark 4.2 for details). If $v$ is a weak solution in the sense of currents to the equation
@@
then $v$ must be constant.
@@ 4.4. Gradient estimates
and by Lemma 2.2, we have

### v3

@@ Abstract.
footnotetext: 2020 Mathematics Subject Classification. Primary 32W20; Secondary 53C55.
@@ 2.1. The $k$-Hessian operators and the special Lagrangian operator
Lemma 2.2 (Yuan [43], Wang-Yuan [41]).
@@ 4.1. Viscosity solutions
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain. Suppose $w$ is an (n1)(n-1)-admissible function and a viscosity subsolution of (45) in $\Omega$, and $v$ is a viscosity supersolution of (45) in $\Omega$. If $w\leq v$ on $\partial\Omega$, then $w\leq v$ in $\Omega$.
@@
Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible. As $v$ is a viscosity supersolution, by the definition of the supersolution, we have
@@ 4.2. Weak solutions in the sense of currents
    If $0\leq v\leq 1$, $\frac{1}{2}v^{2}$ is a subsolution to the same equation.
@@
Let $v$ be a weak subsolution in the sense of currents to the equation $\sigma_{n}+\sigma_{n-1}=0$. Then $v$ is also a viscosity subsolution.
@@ 4.3. A Liouville theorem to the equation $\sigma_{n}+\sigma_{n-1}=0$ in $\mathbb{C}^{n}$
Let $v:\mathbb{C}^{n}\rightarrow\mathbb{R}$ be a $C^{1}$ function that is (n1)(n-1)-subharmonic and such that $v+|z|^{2}$ is plurisubharmonic. Assume there exists a positive constant $C$ such that $\|v\|_{C^{1}(\mathbb{C}^{n})}\leq C$ and $\Delta v\leq C$ in the weak sense (see Remark 4.2 for details). If $v$ is a weak solution in the sense of currents and a viscosity supersolution to the equation
@@
then $v$ must be a constant.
@@ 4.4. Gradient estimates
where the matrix $A^{i}$ is uniformly bounded and by Lemma 2.2, we have
@@
Next, we prove $u^{\infty}$ is a supersolution to the equation $\sigma_{n-1}+a_{0}\sigma_{n}=0$ in the viscosity sense. For any $C^{2}$ function $\varphi$ with $\lambda(\varphi_{\alpha\bar{\beta}})\in\overline{\Gamma}_{n-1}$ and for any $z_{0}$ of the local maximum point of $\varphi-u^{\infty}$, we need to show
$\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(z_{0})\big)+a_{0}\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(z_{0})\big)\leq 0.$ (66)
If $\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))=0$, then $\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\leq 0$ and thus (66) holds.
So we assume $c_{0}=\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))>0$. Then for any $\epsilon>0$ small enough, we have
$\lambda(\varphi_{\alpha\bar{\beta}}-2\epsilon\delta_{\alpha\beta})\in\Gamma_{n-1}\ \ \text{and}\ \ \sigma_{n-1}(\varphi_{\alpha\bar{\beta}}-2\epsilon\delta_{\alpha\beta})>\frac{c_{0}}{2}\ \text{in}\ \overline{B}_{\epsilon}(z_{0}).$ (67)
For sufficiently large $i$ such that $|\hat{u}^{t_{i}}-u^{\infty}|<\frac{\epsilon^{3}}{2}$, consider the function
$\Phi=\varphi-\epsilon|z-z_{0}|^{2}-\hat{u}^{t_{i}}.$
Since for any $z\in\partial B_{\epsilon}(z_{0})$,
$\Phi(z)=\varphi(z)-\hat{u}^{t_{i}}(z)-\epsilon^{3}<$ $\varphi(z)-u^{\infty}(z)-\frac{\epsilon^{3}}{2}$
$\leq$ $\varphi(z_{0})-u^{\infty}(z_{0})-\frac{\epsilon^{3}}{2}<\varphi(z_{0})-\hat{u}^{t_{i}}(z_{0})=\Phi(z_{0}).$
then we know that $\Phi$ attains its maximum on $\overline{B}_{\epsilon}(z_{0})$ at an interior point $\tilde{z}_{0}\in B_{\epsilon}(z_{0})$. Then $(\Phi_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}\leq 0$, namely $(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-\epsilon\delta_{\alpha\beta})_{n\times n}\leq(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}$. Thus $(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}\geq(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta})_{n\times n}$ and we obtain
$\sigma_{n-1}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A_{\alpha\bar{\beta}}(\tilde{z}_{0}))\geq\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)\geq\frac{c_{0}}{2}.$
By the concavity of $\frac{\sigma_{n}}{\sigma_{n-1}}$ in $\Gamma_{n-1}$, we get
$1-\tan\hat{\theta}(t_{i})M^{2}_{i}\frac{\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}{\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}$
$\leq$ $1-\tan\hat{\theta}(t_{i})M^{2}_{i}\frac{\sigma_{n}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))}{\sigma_{n-1}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))}$
$\leq$ $c_{0}^{-1}CM_{i}^{-2},$ (68)
where in the last inequality we use (64). Then we get
${\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}-\tan\hat{\theta}(t_{i})M^{2}_{i}{\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}\leq c_{0}^{-1}CM_{i}^{-2}\sup|D^{2}\varphi|.$ (69)
Let $i\rightarrow\inftyand$ then $\epsilon\rightarrow 0$ in the above, we obtain (66). Hence $u^{\infty}$ is a supersolution to the equation $\sigma_{n}+\sigma_{n-1}=0$.